OpenStudy (anonymous):

My sister has an algebra problem but I haven't done it in so long I can't even answer it... v + 3/5 = 2 1/3

OpenStudy (phi):

this is an equation. subtract 3/5 from both sides: v + 3/5 - 3/5 = 2 1/3 - 3/5 v= 2 1/3 - 3/5 now you need how to do fractions

OpenStudy (phi):

convert 2 1/3 to an improper fraction. 2 is the same as 6/3 so 2 1/3 = 7/3 7/3 - 3/5 use a common denominator of 15